# GRPose: Learning Graph Relations for Human Image Generation with Pose Priors

### v1

@@ Abstract
Recent methods using diffusion models have made significant progress in human image generation with various additional controls such as pose priors. However, existing approaches still struggle to generate high-quality images with consistent pose alignment, resulting in unsatisfactory outputs. In this paper, we propose a framework delving into the graph relations of pose priors to provide control information for human image generation. The main idea is to establish a graph topological structure between the pose priors and latent representation of diffusion models to capture the intrinsic associations between different pose parts. A Progressive Graph Integrator (PGI) is designed to learn the spatial relationships of the pose priors with the graph structure, adopting a hierarchical strategy within an Adapter to gradually propagate information across different pose parts. A pose perception loss is further introduced based on a pretrained pose estimation network to minimize the pose differences. Extensive qualitative and quantitative experiments conducted on the Human-Art and LAION-Human datasets demonstrate that our model achieves superior performance, with a 9.98% increase in pose average precision compared to the latest benchmark model. The code is released on *******.
@@
Figure 1: GRPose generates results by better aligning with pose prior and scaling outputs to $512\times 512$ pixels. The first row compares it with ControlNet while the second row visualizes the pose alignment across different base models. ControlNet exhibits pose alignment in certain body parts such as arms and incorrect actions like handshakes.
**footnotetext: Xun Yang is the corresponding author.
Introduction
@@
Human image generation aims to synthesize high-quality images under specific conditions based on a series of prompts, such as canny edge, pose and depth (Yang et al. 2023). Its diverse applications range from animation (Corona et al. 2024) and game production (Pan et al. 2024) to other fields. Early methods (Men et al. 2020; Ma et al. 2017) primarily adopted variational autoencoders (VAEs) (Kingma and Welling 2013) or Generative Adversarial Networks (GANs) (Goodfellow et al. 2020), leveraging a source image to synthesize target images with specific human attributes. While these methods achieve control through the reference appearance, the synthesis process is unstable and the training heavily depends on the distribution of the source images. Recently, Stable Diffusion (SD) (Rombach et al. 2022) and its variants (Podell et al. 2023) have been developed to address these limitations, in which high-fidelity human images are synthesized with the help of a prompt. Considering the data availability and computation costs, controllable diffusion models (Zhang, Rao, and Agrawala 2023; Mou et al. 2024; Zhao et al. 2024) further introduce a learnable control branch into the frozen SD model, enabling spatial control of the generative results by providing conditions such as depth maps and segmentation masks.
@@
ControlNet (Zhang, Rao, and Agrawala 2023) has emerged as a new benchmark by fine-tuning the frozen SD model with trainable copy parameters, enabling spatial control through conditional inputs. HumanSD (Ju et al. 2023b) adopts a heat-guided loss to achieve pose control. While these methods synthesize images with satisfactory semantics and style, they struggle to produce high-quality outputs based on given pose priors, often resulting in unrealistic body alignment (as shown in Figure 1). A key challenge in existing approaches is the poor alignment with the given pose priors. We observed that these methods incorporate human pose information into frozen generative models based on Euclidean space, which inadequately represents the nonlinear and higher order relationships between different parts of pose priors, particularly in terms of joint connections and overall coordination.
@@
In this study, we propose to capture the topological relationships among different pose parts by leveraging a graph structure to extract intrinsic associations. We propose a framework named Graph Relation Pose (GRPose) to guide Stable Diffusion. We follow the ControlNet utilizing both a text prompt and a pose prior to generate human images, and primarily fine-tune the Adapter which is a set of trainable copy parameters. We aim to establish a graph topological structure on pose priors and latent representations, adopting Graph Convolutional Networks (GCNs) (Kipf and Welling 2016) to effectively capture the higher order relationships between different pose parts. We introduce a graph construction mechanism called the Progressive Graph Integrator (PGI). PGI treats each spatial part of pose priors and latent representations as graph nodes, mapping the spatial relationships of the pose into a graph structure. This mechanism distills pose information into each layer within the Adapter through a hierarchical structure, allowing gradual refinement of the pose while preserving the authenticity of the appearance. Additionally, we propose a novel pose perception loss, which adopts a pre-trained pose estimation network to minimize the pose differences between outputs and original images, further encouraging alignment of the synthesized output with pose priors.
Our contributions are summarized as follows:
@@
  • We propose a framework called Graph Relation Pose (GRPose), leveraging a graph structure to more comprehensively capture the topological and spatial information of pose priors for human image generation, offering a new perspective in the field.
@@
  • We design a Progressive Graph Integrator (PGI) to explicitly capture the intrinsic associations between different pose parts, distilling pose information into each layer within the Adapter.
@@
  • We introduce a novel pose perception loss that utilizes a pre-trained pose estimation network to minimize pose differences.
@@
  • We conduct extensive qualitative and quantitative experiments on the Human-Art (Ju et al. 2023a) and LAION-Human datasets (Ju et al. 2023c). Our GRPose achieves superior performance compared to advanced methods across multiple evaluation metrics, particularly in terms of pose guidance alignment.
@@
Figure 2: Overview of Graph Relation Pose (GRPose). It consists of Diffusion Pipeline, Graph Pose Adapter and Pose Perception Loss. The Pose Encoder uses a coupling structure in alongside the Progressive Graph Integrator (PGI) to capture graph relationships between different pose parts. The Pose Perception Loss adopts a pre-trained pose estimation network to regularize the pose alignment in the feature space.
Related Work
Pose-Guided Human Image Generation
@@
Traditional pose-guided human generation methods take a specific source image and pose condition as input to generate images that retain the appearance of the source image while presenting a specified pose. Some studies (Lv et al. 2021; Ma et al. 2021; Yang et al. 2021) are based on GANs or VAEs to convert the task as a conditional generation. Bhunia et al.  (Bhunia et al. 2023) leveraged a Person Image Diffusion Model (PIDM) to synthesize images by learning a noise distribution and proposed a cross-attention-based Texture Diffusion Module (TDB) to align the relationship between appearance and pose information. Zhang et al.  (Zhang et al. 2022) proposed a dual-task pose transformer and introduced an auxiliary task, connecting different branches to obtain correlation between features by building attitude conversion modules. Shen et al.  (Shen et al. 2023) introduced a Progressive Conditional Diffusion Model (PCDM), which incrementally bridges the gap between character images under target poses and source poses through a three-stage process. Despite the significant progress of existing methods, challenges remain in handling complex poses. Our study seeks to achieve more realistic human images by introducing graph associations and a pose perception loss.
Controllable T2I Diffusion Model
@@
Large Diffusion models (Rombach et al. 2022; Ramesh et al. 2022) can generate high-quality images under a set of prompts. Ho et al.  (Ho and Salimans 2022) proposed classifier-free guidance that combines conditional prediction with unconditional prediction. ControlNet (Zhang, Rao, and Agrawala 2023) introduces a trainable control branch by incorporating a weight copy of the diffusion model, allowing spatial control of generative images using depth maps, segmentation masks and more. HumanSD (Ju et al. 2023b) finetunes the diffusion model with a pose-guided heatmap loss. T2I-Adapter (Mou et al. 2024) trains a lightweight adapter to guide the frozen SD model. UniControlNet (Zhao et al. 2024) utilizes two additional control modules to achieve control over multiple conditions. Although these methods demonstrate strong control capabilities, they struggle with handling complex pose relationships. To address this, we introduce graph structures to capture high-level associations in the pose-guided generation process.
Method
Overview of GRPose
@@
where $z_{t}$ represents the latent representation at time step $t$, $\mathcal{N}$ represents the standard normal distribution, $\epsilon_{\theta}$ represents the noise prediction network and $\epsilon_{t}$ denotes real noise.
Within the entire framework, the Graph Pose Adapter is a trainable component that encodes the pose condition $c_{p}$ into a graph structure and integrates it into the Adapter through a hierarchical structure. At the beginning of each encoder layer in the Adapter, the encoded pose and the current latent representation are fed into the Progressive Graph Integrator (PGI) to capture the topological relationships between different pose parts through graph learning. This process fine-tunes the Adapter to convey control signals to the SD model, producing the synthesized image $\hat{x}$. Additionally, to further encourage alignment of the synthesized output with pose priors, the pose perception loss is formulated using a pre-trained pose estimation network to quantify the pose differences between outputs and the original images.
Progressive Graph Integrator
We adopt a pose encoder with $L$ layers ($L$=4, similar to the Adpater) to capture multi-level scales of pose priors within a hierarchical structure, where a PGI is incorporated into each encoding layer and its corresponding layer in the Adapter. In the PGI, each spatial point in the encoded pose prior $x_{p}$ and the latent representation $x_{l}\in\mathbb{R}^{H\times W\times C}$ is treated as a node for constructing respective graph structures, as shown in Figure 3. By doing this, both local features of spatial relationships in the pose prior and latent representation are captured, where each point in $x_{p}$ or $x_{l}$ represents a field-of-view of $(2^{L-1}\times 2^{L-1})$ in the original images. Then, we proceed to construct the feature graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the node set is defined as $\mathcal{V}=\{v_{1},v_{2},\dots,v_{N}\}$ with $N=H\times W$ representing the number of nodes. The graph edges between nodes $v_{i}$ and $v_{j}$ are established by applying a K-nearest neighbors ($K$NN) search algorithm:
@@
where $d$ represents the $L_{2}$ distance between two node vectors. Notably, we incorporate positional encoding $p_{i}$ into each node ( $v_{i}\leftarrow v_{i}+p_{i}$), where $p_{i}$ denotes the spatial positions of the corresponding patch. A set of edge $\mathcal{E}$ can be obtained as:
@@
Then, two separate graph structures, $\mathcal{G}_{p}$ and $\mathcal{G}_{l}$, are formed to represent the pose prior $x_{p}$ and latent representastion $x_{l}$ respectively.
With these two graphs, we employ graph convolution (GC) layers to facilitate the association of features between nodes within each graph. The GC layer is divided into two steps: aggregation and update. In the aggregation step, features of neighboring nodes are gathered as described:
@@
where $\mathcal{K}(i)$ is the neighboring node $i$ determined by the $K$NN algorithm, $\mathcal{A}\in\mathbb{R}^{N\times N}$ is an adjacency matrix, and $X$ denotes the features of nodes. In the update step, each node integrates the aggregation features to update its current representation. This process is expressed as:
@@
where $\boldsymbol{\phi}$ denotes the activation function, $\boldsymbol{\Theta}$ is a learnable parameter. To balance the degrees of different nodes, the adjacency matrix $\mathcal{A}$ is normalized as:
@@
where $D$ denotes the diagonal degree of $\mathcal{A}$, $I$ denotes the identity matrix. The GC process can be described as:
@@
where $W$ denotes the weight matrix. The GC layer distills information across different parts for both pose prior and latent representation. Take the pose prior as the example, this process is summarized as:
@@
We further fuse the graph features of pose and latent through a fuse layer, which captures the cross-modal interactions between them. The fuse layer consists of three convolution blocks. The fused features are then used to construct a graph $\mathcal{G}_{fuse}$, which is fed into an additional GC layer to capture the complex associations to refine the feature representation.
@@
Pose Perception Loss
We introduce a novel Pose Perception Loss to enhance the pose alignment during the image synthesis process under the guidance of a pose perception network. Our approach employs a pre-trained pose estimation network, specifically designed to accurately identify and comprehend the poses of individuals within images. By encoding both the generated image $\hat{x}$ and the original image $x$, the discrepancies between them can be captured effectively, the it can be defined as:
@@
where $\varphi_{p}$ denotes the encoder of the pose estimation network, which extract the features from the pose backbone, $h$ and $w$ denotes the width and height of both $x$ and $\hat{x}$ respectively and $\left\|\cdot\right\|_{2}$ signifies the Euclidean distance. By directly minimizing the pose differences in the feature space, our method is capable of generating images that are not only visually more realistic but also consistent with pose priors. This loss function can be integrated as a plug-and-play component into existing image generation frameworks, providing additional pose guidance to these models.
The optimization target of the training process is:
@@
where $\mathcal{L}_{d}$ represents the diffusion loss, $\alpha$ respresents the weight of the pose perception loss $\mathcal{L}_{p}$. The parameters of the Adapter are updated through this loss function.
Table 1: Results on Human-Art and LAION-Human datasets. The best results and the second best results are marked in green and blue respectively. Results marked with asterisk (*) are evaluated on the released models.
@@
Experiments
Experimental Settings

### v2

@@ Abstract
Recent methods using diffusion models have made significant progress in human image generation with various control signals such as pose priors. However, existing efforts are still struggling to generate high-quality images with consistent pose alignment, resulting in unsatisfactory output. In this paper, we propose a framework that delves into the graph relations of pose priors to provide control information for human image generation. The main idea is to establish a graph topological structure between the pose priors and latent representation of diffusion models to capture the intrinsic associations between different pose parts. A Progressive Graph Integrator (PGI) is designed to learn the spatial relationships of the pose priors with the graph structure, adopting a hierarchical strategy within an Adapter to gradually propagate information across different pose parts. Besides, a pose perception loss is introduced based on a pretrained pose estimation network to minimize the pose differences. Extensive qualitative and quantitative experiments conducted on the Human-Art and LAION-Human datasets clearly demonstrate that our model can achieve significant performance improvement over the latest benchmark models. The code is available at https://xiangchenyin.github.io/GRPose/.
**footnotetext: Wei Chen and Xun Yang are co-corresponding authors. The work was done within intern at Li Auto.
1 Introduction
@@
Figure 1: Examples of the pose-guided human image generation task. The first row illustrates the generated results of the ControlNet and our GRPose methods while the second row visualizes the pose alignment across different base models. GRPose generates better results by well aligning with pose prior and scaling outputs to $512\times 512$ pixels.
@@
Human image generation aims to synthesize high-quality images under specific conditions based on a series of prompts, such as canny edge, pose and depth (Yang et al. 2023; Wang et al. 2024a). Its diverse applications range from animation (Corona et al. 2024) and game production (Pan et al. 2024) to other fields. Early methods (Men et al. 2020; Ma et al. 2017) primarily adopted variational autoencoders (VAEs) (Kingma and Welling 2013) or Generative Adversarial Networks (GANs) (Goodfellow et al. 2020), leveraging a source image to synthesize target images with specific human attributes. Although these methods achieve control through the reference appearance, the synthesis process is unstable and the training heavily depends on the distribution of the source images. Recently, Stable Diffusion (SD) (Rombach et al. 2022) and its variants (Podell et al. 2023) have been developed to address these limitations, in which high-fidelity human images are synthesized with the help of a prompt. Considering the data availability and computation costs, controllable diffusion models (Zhang, Rao, and Agrawala 2023; Mou et al. 2024) further introduce a learnable control branch into the frozen SD model, enabling spatial control of the generative results based on the provided conditions, e.g., depth maps and segmentation masks.
@@
In recent years, ControlNet (Zhang et al. 2023) has emerged as a new benchmark by fine-tuning the frozen SD model with trainable copy parameters, enabling spatial control through conditional inputs. HumanSD (Ju et al. 2023b) adopts a heat-guided loss to achieve pose control. While these methods synthesize images with satisfactory semantics and style, they are still struggling to produce high-quality output based on the pose priors, often resulting in unrealistic body alignment (as shown in Figure 1). A key challenge in existing approaches is the poor alignment with the given pose priors. We observed that these methods incorporate human pose information into frozen generative models based on Euclidean space, which inadequately models the nonlinear and higher-order relationships between different parts of pose priors, particularly in terms of joint connections and overall coordination.
@@
In this study, we propose to capture the topological relationships among different pose parts by leveraging a graph structure to extract intrinsic associations. We propose a framework named Graph Relation Pose (GRPose) to guide the process of stable diffusion. We follow the ControlNet framework that utilizes both a text prompt and a pose prior to generate human images, and primarily fine-tune the Adapter which is a set of trainable copy parameters. We aim to establish a graph topological structure on pose priors and latent representations, adopting Graph Convolutional Networks (GCNs) (Kipf et al. 2016; Han et al. 2022) to effectively capture the higher-order relationships between different pose parts. We introduce a graph construction mechanism called the Progressive Graph Integrator (PGI). PGI treats each spatial part of pose priors and latent representations as graph nodes, mapping the spatial relationships of the pose into a graph structure. This mechanism distills pose information into each layer within the Adapter through a hierarchical structure, allowing gradual refinement of the pose while preserving the authenticity of the appearance. Additionally, we propose a novel pose perception loss, which adopts a pretrained pose estimation network to minimize the pose differences between outputs and original images, further encouraging alignment of the synthesized output with pose priors. Our contributions are summarized as follows:
@@
  • We propose a novel framework, termed GRPose, that exploits a graph structure to effectively capture the topological and spatial information of pose priors for human image generation, offering a new perspective in the field.
@@
  • We design a Progressive Graph Integrator (PGI) to explicitly capture the intrinsic associations between different pose parts, distilling pose information into each layer.
@@
  • We introduce a novel pose perception loss that utilizes a pose estimation network to minimize pose differences.
@@
  • We conduct extensive experiments on the Human-Art (Ju et al. 2023a) and LAION-Human datasets (Ju et al. 2023c). Our GRPose achieves superior performance compared with advanced methods across multiple metrics, particularly in terms of pose guidance alignment.
@@
Figure 2: Overview of Graph Relation Pose (GRPose). The Pose Encoder is adopted to capture multi-level scales of pose priors within a hierarchical structure, where a Progressive Graph Integrator is incorporated to capture graph relationships between different pose parts. The Pose Perception Loss adopts a pre-trained pose estimation network to regularize the pose alignment.
2 Related Work
2.1 Pose-Guided Human Image Generation
@@
Traditional pose-guided human generation methods take a specific source image and pose condition as input to generate images that retain the appearance of the source image while presenting a specified pose. Recently, deep learning techniques have achieved significant progress in various domains (Fan et al. 2022; Liang et al. 2024). Some studies (Lv et al. 2021; Ma et al. 2021; Yang et al. 2021) are based on GANs or VAEs to convert the task as a conditional generation. Bhunia et al.  (Bhunia et al. 2023) leveraged a Person Image Diffusion Model (PIDM) to synthesize images by learning a noise distribution and proposed a cross-attention-based Texture Diffusion Module (TDB) to align the relationship between appearance and pose information. Zhang et al.  (Zhang et al. 2022) proposed a dual-task pose transformer and introduced an auxiliary task, connecting different branches to obtain correlation between features by building attitude conversion modules. Shen et al.  (Shen et al. 2023) introduced a Progressive Conditional Diffusion Model (PCDM), which incrementally bridges the gap between character images under target poses and source poses through a three-stage process. HumanSD (Ju et al. 2023b) finetuned the diffusion model with a pose-guided heatmap loss and created a new LAION-Human dataset. Stable-Pose (Wang et al. 2024b) introduced a coarse-to-fine attention masking strategy into ViT to obtain more accurate pose guidance. Despite the significant progress of existing methods, challenges remain in handling complex poses.
2.2 Controllable T2I Diffusion Model
@@
Large Diffusion models (Rombach et al. 2022; Ramesh et al. 2022) generate high-quality images under a set of prompts. Ho et al.  (Ho and Salimans 2022) proposed classifier-free guidance that combines conditional prediction with unconditional prediction. ControlNet (Zhang, Rao, and Agrawala 2023) introduces a trainable control branch by incorporating a weight copy of the diffusion model, allowing spatial control of generative images using depth maps, segmentation masks and more. T2I-Adapter (Mou et al. 2024) trains a lightweight adapter to guide the frozen SD model. UniControlNet (Zhao et al. 2024) utilizes two additional control modules to achieve control over multiple conditions. Although these methods demonstrate strong control capabilities, they struggle with handling complex pose relationships.
3 Methodology
3.1 Overview of Our GRPose
@@
where $z_{t}$ denotes the latent representation at time step $t$, $\mathcal{N}$ denotes the standard normal distribution, $\epsilon_{\theta}$ denotes the noise prediction network and $\epsilon_{t}$ denotes real noise.
Within the entire framework, the Graph Pose Adapter is a trainable component that encodes the pose condition $c_{p}$ into a graph structure and integrates it into the Adapter through a hierarchical structure. At the beginning of each encoder layer in the Adapter, the encoded pose and the current latent representation are fed into the Progressive Graph Integrator (PGI) to capture the topological relationships between different pose parts through graph learning. This process fine-tunes the Adapter to transfer the control signals into the SD model, producing the synthesized image $\hat{x}$. Additionally, to further encourage the alignment of synthesized output with pose priors, the pose perception loss is formulated using a pretrained pose estimation network to quantify the pose differences between the outputs and original images.
3.2 Progressive Graph Integrator
We adopt a pose encoder with $L$ layers ($L$=4, similar to the Adpater) to capture multi-level scales of pose priors within a hierarchical structure, where a PGI is incorporated into each encoding layer and its corresponding layer in the Adapter. In the PGI, each spatial point in the encoded pose prior $x_{p}$ and the latent representation $x_{l}\in\mathbb{R}^{H\times W\times C}$ is treated as a node to construct the graph structures, as shown in Figure 3. By doing this, both local features of spatial relationships in the pose prior and latent representation are captured. Then, we proceed to construct the feature graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the node set is defined as $\mathcal{V}=\{v_{1},v_{2},\dots,v_{N}\}$ where $N=H\times W$ denotes the number of nodes. The graph edges between nodes $v_{i}$ and $v_{j}$ are established by applying a K-nearest neighbors ($K$NN) search algorithm:
@@ 3.2 Progressive Graph Integrator
where $d$ denotes the $L_{2}$ distance between two node vectors. Notably, we incorporate positional encoding $p_{i}$ into each node ( $v_{i}\leftarrow v_{i}+p_{i}$), where $p_{i}$ denotes the spatial positions of the corresponding patch. A set of edge $\mathcal{E}$ can be obtained as:
@@
Then, two separate graph structures, $\mathcal{G}_{p}$ and $\mathcal{G}_{l}$, are formed to represent the pose prior $x_{p}$ and latent representastion $x_{l}$ respectively. With these two graphs, we employ graph convolution (GC) layers to facilitate the association of features between nodes within each graph. The GC layer is divided into two steps: aggregation and update. In the aggregation step, features of neighboring nodes are gathered as follows:
@@
where $\mathcal{K}(i)$ denotes the set of the neighboring nodes of the node $i$ determined by the $K$NN algorithm, $\mathcal{A}\in\mathbb{R}^{N\times N}$ is an adjacency matrix, and $X$ denotes the features of nodes. In the update step, the current node representation is updated based on the aggregated features, as follows:
@@
where $\boldsymbol{\phi}$ denotes the activation function, $\boldsymbol{\Theta}$ is a learnable parameter. Then, the adjacency matrix $\mathcal{A}$ is normalized as:
@@
where $D$ denotes the diagonal degree of $\mathcal{A}$, and $I$ denotes the identity matrix. The GC process can be described as:
@@
where $W$ denotes the weight matrix. The GC layer distills information across different parts for both pose prior and latent representation. Take the pose prior as an example, this process is summarized as:
@@
We further fuse the graph features of pose and latent through a fusion layer, which captures the cross-modal interactions. The fusion layer consists of three convolution blocks. The fused features are then used to construct a graph $\mathcal{G}_{fuse}$, which is fed into an additional GC layer to capture the complex associations to refine the feature representation.
@@
3.3 Pose Perception Loss
We introduce a novel Pose Perception Loss to enhance the pose alignment during the image synthesis process under the guidance of a pose perception network. Our approach employs a pre-trained pose estimation network, specifically designed to accurately identify and comprehend the poses of individuals within images. By encoding both the generated image $\hat{x}$ and the original image $x$, the discrepancies between them can be captured effectively, defined as follows:
@@ 3.3 Pose Perception Loss
where $\varphi_{p}$ denotes the encoder of the pose estimation network, which extract the features from the pose backbone, $h$ and $w$ denotes the width and height of both $x$ and $\hat{x}$ respectively and $\left\|\cdot\right\|_{2}$ denotes the Euclidean distance. By directly minimizing the pose differences in the feature space, our method is capable of generating images that are not only visually more realistic but also consistent with pose priors. This loss function can be integrated as a plug-and-play component into existing image generation frameworks, providing additional pose guidance to these models. The optimization objective of the training process is formulated as:
@@
where $\mathcal{L}_{d}$ denotes the diffusion loss, and $\alpha$ denotes the weight of the pose perception loss $\mathcal{L}_{p}$. The parameters of the Adapter are updated through this loss function.
Table 1: Results on the Human-Art and LAION-Human datasets. The best results and the second best results are marked in green and blue, respectively. Results marked with asterisk (*) are reimplemented based on the released models.
@@
4 Experiments
4.1 Experimental Settings